A circular conducting loop of radius 23.0 cm is located in a region of homogeneous magnetic field of magnitude 0.500 T pointing perpendicular to the plane of the loop. The loop is connected in series with a resistor of 167 Ω. The magnetic field is now increased at a constant rate by a factor of 2.20 in 19.0s.
1.Calculate the magnitude of the induced emf in the loop while the magnetic field is increasing.
2.Calculate the magnitude of the current induced in the loop while the field is increasing.
3.With the magnetic field held constant at its new value of 1.10 T, calculate the magnitude of the average induced voltage in the loop while it is pulled horizontally out of the magnetic field region during a time interval of 7.50 s.
Answer #1

given

dB/dt = (2.20-0.5)/19 = 0.089 T/s

Area of loop = A = \pi r2 = \pi (0.23)2 = 0.166 m2

1) induce emf = E = A dB/dt

E = 0.166 x 0.089 = 0.014 volts

2) as resistance =R = 167 ohm

so so current i = E/R = 0.014 / 167 = 0.088 mA

3) now dB/dt = (1.10 -0.5)/7.5 = 0.08 T/s

so emf = A dB/dt =  0.166 x 0.08 = 0.013 volts

so current i = 0.013 / 167 = 0.079 mA

answer